The circle (x – 4)2 + (y – 3)2 = 4 has centre C and radius r.
The line y = x + 1 intersects this circle at two points A and B.
Find the coordinates of A and B.
Some working out would be appreciated to learn for next time :D
Thanks in advance

Answer by stanbon(75887)

The circle (x – 4)2 + (y – 3)2 = 4 has centre C and radius r.
The line y = x + 1 intersects this circle at two points A and B.
Find the coordinates of A and B.
------
Substitute for "y" and solve for "x":
(x-4)^2 + (x+1-3)^2 = 4
---
x^2 -8x +16 + (x^2-4x+4) = 4
----
2x^2 -12x +16 = 0
---
x^2 -6x +8 = 0
Factor:
(x-4)(x-2) = 0
x = 2 or x = 4
----
Solve for the corresponding "y": values using y = x+1
If x = 2, y = 3
If x = 4, y = 5
------
Points of intersection: (2,3) and (4,5)
=======================================
Cheers,
Stan H.